OpenStudy (anonymous):

if sin (theta) = x/3 then sin (2 theta) = 2x/3 is this true?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

not always

OpenStudy (anonymous):

\[\sin \theta = \frac{ x }{ 3 }\] \[\sin 2\theta = \frac{ 2x }{ 3 }\] if no, then what is \[\sin 2\theta \]

OpenStudy (amoodarya):

\[\sin 2\theta=2 \sin \theta \cos \theta\]

OpenStudy (amoodarya):

\[\sin ^2 \theta +\cos ^2 \theta=1 \\cos^2 \theta=1-\sin ^2 \theta\\cos \theta =\pm \sqrt{1-\sin^2 \theta}\]

OpenStudy (anonymous):

the original question was \[\int\limits_{}^{} \frac{ 8 dx }{ (4x ^{2} +1)^{2} } \]

OpenStudy (amoodarya):

use substitution x=1/2 tan u dx=1/2 (1+tan ^2 u+1) du